OpenStudy (anonymous):

(cos Θ + cos Θ)2 + (cos Θ + cos Θ)2 Answer options: A. sin2 Θ B. 4 C. 8cos2 Θ D. cos2 Θ

OpenStudy (anonymous):

@ganeshie8 @hartnn

OpenStudy (cwrw238):

(cos Θ + cos Θ)2 = (2cosΘ)^2 = 4cos^2Θ

OpenStudy (gorv):

\[(2\cos \theta)^2+(2\cos \theta)^2=4\cos^2+4\cos^2=8\cos^2 \theta\]